OpenStudy (anonymous):

what is the square root of 306?

OpenStudy (anonymous):

sqrt(306) = 17.4928557

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

no that was it. I was working on the pythagorean theorm.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Random is good... random is good hehe

OpenStudy (anonymous):

\[\sqrt{306}=\sqrt{17*18}=\sqrt{17*2*9}=3\sqrt{34}\]